OpenStudy (anonymous):

find tan theta if cos theta = 1/4 and sin theta < 0 use the fundamental identities to find the value of the trigonometric function

OpenStudy (goformit100):

For solving this do you know the conception of : Trigonometric Functions.

OpenStudy (anonymous):

YES can you actually help me get the answer tho?

OpenStudy (anonymous):

ok leave that. tell me, how do you relate sec theta with cos theta?

OpenStudy (anonymous):

sec theta = 1/cos theta

OpenStudy (anonymous):

yes. so sec theta = 1/1/4 = 4. Now which identity relates sec theta with tan theta?>

OpenStudy (anonymous):

1 + tan^2 theta = sec^2 theta

OpenStudy (anonymous):

yes. So 1 + Tan^2 theta = 4^2 Tan^2 theta = 15. Hence Tan theta = sqrt(15). You never needed sin for that! :P

OpenStudy (jdoe0001):

if \(sin(\theta) <0\) meaning negative, \(tan(\theta)\) will also be negative

OpenStudy (jdoe0001):

meaning either III or IV quadrant

OpenStudy (anonymous):

ok that totally helped me understand, but for some reason the answer options are